first and last gear having teeth 30 and 50 respectively of a simple gear train what will be train value and speed ratio gear respectively if first gear is driving gear


1
Expert's answer
2021-07-02T16:30:02-0400

Train value

"= \frac{Driving \space gear}{Driven \space gear}= \frac{30}{50}= \frac{3}{5}"

Speed ratio gear

"= \frac{Driven \space gear}{Driving \space gear}= \frac{50}{30}= \frac{5}{3}"